Solve for w: w/4 + w/3 = 28
Answer by amoresroy(361)

Solve for w: w/4 + w/3 = 28
You can rewrite the equation as
w(1/4+1/3) = 28
w(3+4)/12 = 28
7w = 12*28
w = 48